Central Corp.s CFO has decided to take a closer look at the firms short-term assets and liabilities. Central Corp.s balance sheet follows Balance Sheet Cash Accounts receivable Inventory Total current assetS $200,000 $110,000 $170,000 $480,000 Accounts payable Accruals Notes payable Total current liabilities Long-term debt Total common equity Total liabilities and equity $170,000 $55,000 $135,000 $360,000 $350,000 $190,000 $900,000 Net plant and equipment $420,000 $900,000 Total assets The value of Central Corp.s working capital is , while its net working capital is The value of Central Corp.s net operating working capital is Central Corp.s current ratio is If Central Corp. decides to purchase a new building with long-term debt, its current ratio will
Answer #1

Working Capital is the total of the current assets of the Company. Hence the total value of Working Capital of Central Corp is $480,000.

Net Working Capital is the difference between the current assets and current liabilities of the Company i.e. Current Assets – Current Liabilities. Hence the Net Working Capital of Central Corp is 480,000 – 360,000 = $120,000.

Net Operating Working Capital is the difference between the operating assets and operating liabilities of the Company i.e. Operating Assets – Operating Liabilities.

Operating assets are those assets acquired for use in the conduct of the regular operations of a business and which are needed to generate revenue for the business. Operating Liabilities are short term liabilities resulting from regular business operations of a Company. These are usually non-interest bearing liabilities.

In the case of Central Corp, all the assets are operating assets the total value of which is $900,000. The operating liabilities are Accounts payable, Accruals and Notes payable. The total operating liabilities are $360,000. Therefore, net operating working capital = 900,000 – 360,000 = $540,000.

Current Ratio = Current Assets / Current liabilities = 480,000/360,000 = 1.33.

If Central Corp decides to purchase a new building with long-term debt, there will be an increase in the Fixed Assets and long term liabilities. Fixed assets are not current assets whereas long term liabilities are not considered as current liabilities. As the current ratio is based on the current assets and current liabilities of a Company, it will not change with the mentioned transaction. The current ratio will still be 1.33.


answered by: ANURANJAN SARSAM